OpenStudy (anonymous):

there are 10 people, 6 are boys and 4 are girls. what is the probability that, when a group of 5 are randomly chosen, that Kevin is chosen? please help, im stuck =(

OpenStudy (anonymous):

let a be kevin, and bcdefghij be the less important kids

OpenStudy (anonymous):

I can pick abcde acdef adefg aefgh afghi aghij so 5 combinations that include kevin

OpenStudy (anonymous):

so its 1 choose 1 times 9 choose 4 dividied by 10 choose 5?

OpenStudy (anonymous):

arent there 6 with kevin?

OpenStudy (anonymous):

I'm mistaken; there are more with kevin

OpenStudy (anonymous):

i think if it is like i said it was, then it should be probability = 1/2 that it includes Mr. Kevin... right?

OpenStudy (anonymous):

thinking too hard

OpenStudy (anonymous):

5 is half of 10. the probability that "kevin" is chosen is \[\frac{1}{2}\]

OpenStudy (anonymous):

yep thats what i got too buddy =) you rule! ...

OpenStudy (anonymous):

cuz its 1/10+1/10+1/10+1/10+1/10 = 5/10... which is one way to get there

OpenStudy (anonymous):

you can use a bunch of formulas if you like, but they cloud the obvious

OpenStudy (anonymous):

you could use \[\frac{\dbinom{1}{1}\times \dbinom{9}{4}}{\dbinom{10}{5}}\]

OpenStudy (anonymous):

or your could use \[1-\frac{9}{10}\times \frac{8}{9}\times\frac{7}{8}\times \frac{6}{7}\times \frac{5}{6}\] that is one minus the probability that kevin is not chosen

OpenStudy (anonymous):

but you are choosing half the people and therefore the probability that kevin is one of them is \[\frac{1}{2}\] without any formulas